Shaypop:

If TX= 2b-27 and VX= b, find the value of b in parallelogram TUVW.

Shaypop:

1 attachment
AZ:

VX = TX

AZ:

set them equal to each other and solve

Shaypop:

Why u said that

snowflake0531:

2b-27=b solve for b

Shaypop:

How do I solve that

snowflake0531:

add 27 to each side

snowflake0531:

and then subtract b

Shaypop:

27

snowflake0531:

YES

snowflake0531:

so, 27

Shaypop:

Don’t u do 27 divided by 2

snowflake0531:

no, since there's only 1 b left

snowflake0531:

2b-27=b b-27 b=27

Shaypop:

In the pic

snowflake0531:

it just says find the value of b